OpenStudy (anonymous):

Interval notation for (x^2+3x)/(x-1) >0

OpenStudy (freckles):

When is x^2+3x=0 and when is x-1=0?

OpenStudy (anonymous):

0 ,-3, and 1

OpenStudy (freckles):

right so (x^2+3x)/(x-1) is zero when x=0 or -3 and is undefined when x=1 but anyways we are going to draw a number line with these numbers on it now and test some intervals

OpenStudy (freckles):